OpenStudy (anonymous):

a girl lifts a 160-N load to a height of 1m in 0.5s. how much power is used to lift the load? 32w 160w 320w 80w

OpenStudy (anonymous):

force applied(F) = 160 N distance (d) = 1m now, woek done by the girl (W)= F * d = 160 * 1m 160 joule we have, time (t) = 0.5 sec therefore, power (P) = w/t =160/0.5 = 320 watt so,320 watt is your correct answer.